I know that the motion is non-uniform, but I am not quite sure how to solve for the average speed.
Measuring Velocity 02 (Active-HTML5) Due this Friday, Jan 25 at 11:59 pm (EST) 10.01s get credit. Remember that you can click and drag the left mouse button to find the location of the cursor time interval between t-4.0s and t-6.0s? (Assume that the mouse coordinates are You need to answer both of these questions correctly to that you can labelled in units of meters. Use the buttons to control the animation. The ghosts are produced at 1 second Intervals.) Click animation and the Start button to start playing the animation. the Initialization button to set up the And second, is the motion uniform or non-uniform? Uniform O Non-Uniform Submit Aw Tries 2/10 Previous Tries This discussion is closed 47 Send Feedback
Answer #1

Ans

The average speed between t = 4s to t = 6s = (distance travelled)/ (time interval)

From the graph, the distance traveled = 5.5 m(( assuming the scale is 1m)

and time interval = 6- 4 = 2 s

The average speed = 5.5/2 = 2.75 m/s

10.01 s